OpenStudy (anonymous):

Can anyone answer this? Last question or at least guide me through. 3 square root of 45 x cubed end root plus 2 square root of 12 end root plus square root of 27 end root minus 3 square root of 20 x cubed 3 x square root of 5 x end root plus 7 square root of 3 3 x square root of 10 x end root plus 7 square root of 6 10 x square root of 8 x 40 x square root of x

OpenStudy (anonymous):

Let me try to re-write it first...

OpenStudy (anonymous):

Sounds good!

OpenStudy (anonymous):

\[3\sqrt{45 x^3} + 2\sqrt{12} + \sqrt{27} - 3\sqrt{20x^3}\]

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

If that problem is written correctly, your calculations are a bit off. Factor everything under each of the square roots.

OpenStudy (anonymous):

Sorry im back and yes thats right

OpenStudy (anonymous):

Just to get your started: \[3 \sqrt{45x^3} = 3 * \sqrt{9} * \sqrt{5} * \sqrt{x^2} * \sqrt{x}\] 9 and \(x^2\) are perfect squares to take the square root and multiply those by the coefficient. \[3 * 3 * x * \sqrt{5} * \sqrt{x}\] Then combine: \[9x\sqrt{5x}\]

OpenStudy (anonymous):

Do the same with the other terms.

OpenStudy (anonymous):

OK thanks i get back to you in a few

OpenStudy (anonymous):

I'll be around.... :-)